# How Haag-tied is QFT, really?

### 18 Dec 2023 v2

@@ 4 The Framework: Assessment, Repair, and Renovation and Maintenance
Agreement (2) belies the different expectations one might have for the form of QFT. Beyond furnishing a method of generating predictions, one pre-Haagian appeal of the interaction picture was its intuitive representation of interaction dynamics. The asymptotic nature of calculations aside, interactions looked familiar: the full state space was time-evolved, from beginning to end, by a specified dynamical operator. As the agreement on (2) reflects, Haag’s theorem undermines the interaction picture’s ability to represent interactions in this way. And while renormalization delivers the agreed-reliable predictions of (1), it does so at the cost of further complicating attempts to represent interactions as originally hoped. But whether any of this is problematic, and if so to what degree, depends on one’s hopes and expectations. What should count as a fully satisfactory theory? as a fully satisfactory theory of relativistic QFT?
In thus organizing the various responses to Haag’s theorem, and in bringing the background commitments of the extra-Haagian outlooks into broad daylight, the framework itself does not adjudicate among the responses to Haag’s theorem. Rather, it clarifies the various lines of genuine debate, both at the level of disagreement of technical matters of fact, and at the level of deeper methodological commitments. As such, these outlooks are directly implicated in 5. Since there is no single “owner” of QFT, let alone the edifice of science, the information that any inquirer of QFT will want—be they physicist, philosopher, or mathematician—will vary according to their own outlook. In our role as ‘general contractors’, we indulge these owners by providing sample advice that either arranges the responses according to their outlook and even provisionally adopts an outlook.
A final note before we proceed. While the responses to Haag’s theorem we review below proceed as if it presents a single problem demanding a singular response, this is not universal. In particular, [Ruetsche, 2011, Chs. 9–11] notes several potential problems that unitary inequivalence phenomena, Haag’s theorem among them, pose to a fundamental particle interpretation. Her response to these is pluralist insofar as she recognizes that “different interpretations will be indexed to—and adulterated by—different aspirations,” where these aspirations “adulterate because they arise along with particular applications of the theory” [Ruetsche, 2011, 246]. She argues that to demand a univocal interpretation, which would require a singular response to Haag’s theorem, would “stym[ie] attempts to harness the theory […] in fruitful projects of explanation and expansion” [Ruetsche, 2011, 260]. While this pluralist response is rich and worth discussion, we do not review it here because of its complexity and because it is not centered on Haag’s theorem per se.
@@
Abandon the interaction picture; adopt Haag-Ruelle scattering
Interpret and assess formal variatns of QFT
Pristine Theory Interpreters
Standard particle notion inconsistent with axioms of QFT
Abandon standard particle interpretation
Develop univocal interpretation of axioms of QFT
@@
Renormalization (breaks Poincaré invariance)
Continue studying renormalization and EFTs, don’t worry
@@
Revise axioms to introduce new evolution parameter
@@
D. Fraser’s dissertation, Haag’s Theorem and the Interpretation of Quantum Field Theories with Interactions, set the stage for much of the philosophical literature on Haag’s theorem over the last sixteen years ([Fraser, 2006]). We will therefore give a more detailed assessment of this source than subsequent works. She writes, “Haag’s theorem frames answers to the question ‘What is QFT?’ ” (p. 171). There are several contenders for the prize of being QFT. Fraser divides what we have called “canonical QFT” into three subcategories: unrenormalized canonical QFT, renormalized canonical QFT, and canonical QFT with cutoffs. Contending against these for the prize of being QFT are several different formal variants of QFT: bottom up axiomatic work based on either the Wightman or Haag-Kastler axioms; constructive QFT; and algebraic QFT. Haag’s theorem shows us that unrenormalized canonical QFT is predicated on an inconsistent set of initial assumptions, and is therefore disqualified. Neither renormalization nor the introduction of cut-offs are satisfactory in the face of this inconsistency:
@@
[C]onsistency is also a relevant criterion because quantum field theory is, by definition, the theory that integrates quantum theory and the special theory of relativity. Consistency is relevant to QFT for theoretical reasons—not for practical reasons (e.g., the derivation of predictions). As a result, it is necessary to either formulate a consistent theory or else show that this criterion cannot be satisfied (i.e., that there is no consistent theory with both quantum and special relativistic principles). …The formal variant is the only variant that satisfies the criterion; its set of theoretical principles are both consistent and well motivated. [Fraser, 2009a] (p. 563)
@@
Thus we have the key dialectical currents of the Faser-Wallace debate fueling this line of response to Haag’s theorem. Bedrock issues about what QFT even is, and what it takes to formulate it in a way that is apt for deep philosophical work on its conceptual-nomological machinery place Haag’s theorem in the spotlight: this no-go theorem regiments answers to the central question at hand, namely, What is QFT? In contrast, the average theoretical particle physicist contents themselves with combining SR and QM to the best of their current ability, rather than demanding perfectly coherent and mathematically rigorous unification up front. At least for now, our current best practice gives up, among other things, the strict Poincaré invariance demanded by SR. Arguably, the ‘best of their ability’ has improved over the decades such that, the average theoretical particle physicist could claim, we will inevitably find such a complete and honest marriage if one is forthcoming. In the meantime, we have successful, if incomplete, marriages of SR with QM with which we can keep experimental and theoretical particle physics going. While it is likewise a live option for D. Fraser (and presumably Earman) that an honest and complete marriage of QFT is not possible, this is a question of logic, not practical or empirical feasibility.141414 Of course, D. Fraser’s full corpus on the philosophy of QFT extends beyond the scope of issues in this present paper. In her dissertation she seemed optimistic that, with more time, the as-of-yet unfinished project of constructing a fully rigorous interacting QFT in four spacetime dimensions could be completed. Crucially, she allowed for the possibility that this will require revision to the axioms ([Fraser, 2006] section 5). Subsequently, she put it to the proponents of renormalized and cutoff variants to wrestle with the question of whether such a completion is impossible: “Neither the infinitely renormalized nor cutoff variant furnishes an argument that a consistent formulation of QFT is impossible; such an argument would require making the case that the axiomatic program cannot be completed.” [Fraser, 2009a] (p. 563). See [Koberinski and Fraser, 2023] for her most recent views on the epistemological status of effective field theory approaches in particle physics.
@@
We note briefly here that the physicists’ attitude of settling for good-enough-to-be-getting-on-with variants of QFT clearly extends far beyond Haag’s theorem. As just one other example, theoretical particle physicists are not bothered when they learn that their QFT path integrals cannot actually be integrals, given that there is no measure on the infinite-dimensional space of paths. Sorting out where path integrals properly live in the mathematical universe is a mathematician’s job—not a physicist’s. Meanwhile, that these path integrals are currently mathematically homeless in no way diminishes their calculational and conceptual fruitfulness for theoretical physics. Haag’s theorem is one result among many indicating that the formal and canonical variants of QFT are not communicating well with each other: either, as Earman and D. Fraser see it, the foundations of QFT are in excellent shape, and they are waiting for more mathematical success in erecting interacting models of realistic matter in a realistic number of spacetime dimensions; or, as Wallace sees it, QFT needs no foundational work since it doesn’t really even need (mathematically) rigorous foundations. We, the authors, think that this communication breakdown between formal and informal variants of QFT is problematic. We will return to these concerns in section 5.
@@
In addition, Earman and D. Fraser recognize that an explanation is needed for the success of the interaction picture and perturbation theory in the face of Haag’s theorem. They “suspect that the full explanation will have a number of different pieces” and offer their analysis of Haag-Ruelle scattering as one such piece (p. 322). The Duncan & Miller response discussed below (section 4.3) should be thought of as adding one additional such piece, not refuting Earman and D. Fraser.
@@
As a consequence of this logical diagnosis and logical remedy, Earman and D. Fraser advocate for future philosophical work and resources to be deployed in philosophical projects about formal variants of QFT and their attendant non-interaction-picture approaches to interactions. There are at least two such broad philosophical projects that benefit from the ways in which practicing mathematicians have “digested the lesson of Haag’s theorem” [Earman and Fraser, 2006] (p. 334). First there is the project of interpreting the mathematical structures used in algebraic QFT and in constructive QFT. Second, and relatedly, there is the subtle matter of assessing the implications of theorems of axiomatic approaches to QFT for the philosophy of those areas of fundamental physics that make use of QFT.
@@ 4.2 Theory Interpreters—The Particle Problem
There is a segment of the philosophical literature addressing Haag’s theorem that is primarily motivated by the question, as put by Ruetsche, ‘Is particle physics particle physics?’ [Ruetsche, 2011, 190]. The issue is whether or not the quantum field theories employed in particle physics admit of particle interpretations. The majority view is that the answer is ‘no’, with Haag’s theorem playing a staring role some lines of reasoning within a larger body of evidence in support of this position. We include an application of our framework to the lines of reasoning within this literature that appeal to Haag’s theorem. However, since many of the arguments at play in the works surveyed below extend beyond the reach of Haag’s theorem, we will not attempt a comprehensive analysis of the arguments for and against a particle interpretation of QFT. [^18] Readers interested in the many lines of argument here beyond those involving Haag’s theorem are encouraged to see chapter 11 of [Ruetsche, 2011] and references therein.
Most notably [Halvorson and Clifton, 2002] and [Fraser, 2008] treat Haag’s theorem as a no-go result for the particle interpretation of QFT. The heart of the argument runs, roughly, as follows. At a minimum, a particle interpretation requires that the theory admits of a formulation in which a number operator exists—you can’t very well say that the theory says it is possible that matter is fundamentally particles if the theory also says that it is not possible (in theoretical principle) for those particles to be counted. The Fock space representation of a free field yields a well-defined number operator suitable for a particle interpretation. Might we then, by means of the IP, use the Fock space of a free field for representing and counting particle states in an interacting system as well? Haag’s theorem blocks this path, since it entails that representations of the ETCCRs for free and interacting systems are unitarily inequivalent. (see Assessment below for more details). Others disagree with this assessment (e.g. [Wallace, 2011] and [Bain, 2011]).
@@
What unifies this sub-literature on Haag’s theorem is that it aims at a decidedly interpretive task: authors of such work aim to answer questions about what ontological commitments are supported by QFT. Thus, this literature is not ultimately motivated by considerations of theory consistency, application, development and innovation, or conceptual-nomological machinery (‘how does this theory work?’) that guide the work of groups surveyed in other subsections. Although such considerations sometime arise in this literature (as, indeed, matters of theory interpretation sometimes arises as a secondary consideration in other segments of the Haag’s theorem literature), they do so as steps within larger arguments ultimately aimed at interpretation. (And of course, many authors take up distinctive projects across this interpretation/not-interpretation divide.)
Within this interpretation-focused literature, the further extra-Haagian outlooks driving the discussion concern the nature of theory interpretation: what is the goal of interpretation; what does it take for a theory to be formulated in a way that is appropriate to the task of interpretation; and to what extent do the contingent details of how the theory is applied in various scientific contexts matter for the project of interpretation. Much of this literature tacitly adheres to what [Ruetsche, 2011] calls the ideal of pristine interpretation. In pristine interpretation, a theory is formulated cleanly in terms of a few powerful laws and principles. The interpreter then seeks to determine where, among all possible worlds, the theory draws a sharp line in “one fell swoop” between what is physically possible and what is physically impossible (p. 3). Contingent details of how the theory gets applied in various contexts are of no consequence for the project of pristine interpretation.
@@
As Ruetsche noted when coining the phrase, this ideal is rarely explicitly stated (p. 3). Nevertheless, sympathy for the ideal of pristine interpretation appears widespread, particularly in the philosophy of QFT. [^19] See [Arageorgis, 1995, 119–125] for an early interpretation-focused discussion of Haag’s theorem. We do not discuss this work here because its status as pristine or not is unclear. [Halvorson and Clifton, 2002] and [Fraser, 2008] approach the particle problem from a methodological stance of pristine interpretation. The pristine interpreters within this group are engaged with a project that has certain commitments from the outset regarding what it takes for a physical theory to be ripe for interpretation. On this approach, the theory to be interpreted is the formal, mathematically well-posed version of the scientific achievement under study ([Fraser, 2009a], [Fraser, 2011]). No-go theorems proved within these formal theoretical structures, then, powerfully and persuasively constrain interpretive options  [^20] See [Freeborn et al., 2023] for further discussion on alternative stances toward no-go theorems in general, and Haag’s theorem in particular. [Fraser, 2008] is explicit on this point. In describing the role of Haag’s theorem for blocking one attempt at developing a particle interpretation of interaction QFT, she writes:
@@
Moreover, according to Wightman, it is a “necessary condition” that any physically sensible interacting field theory use a representation of the ETCCRs that is unitarily equivalent to a Fock representation. Having found a necessary condition for a particle interpretation, pristine interpreters use this as the clean dividing line between theories that do admit of a particle interpretation (i.e. free fields) and those that do not (i.e. interacting fields). Thus, a presupposition of favoring mathematical rigor and axiomatic approaches is at play in the background of these pristine interpretations of Haag’s theorem. [^21] See [Hans Halvorson, 2007] p. 731-732 for a concise statement of this attitude towards interpretation and mathematical precision in theory formulation.
Non-pristine interpreters, in contrast, take a more flexible and opportunistic approach to no-go theorems. To them, Wightman’s “necessary condition” looks more like an interesting suggestion or a useful starting point. From such a starting point, they turn their attention toward opportunities within the messy details of how interacting QFT gets applied to develop serviceable notion of particles without meeting the stringent requirements of finding a representation unitarily equivalent to a Fock space representation. The usual response along these lines appeals to a scale-variant or emergent particle interpretation ([Wallace, 2011, 124], [Wallace, 2001]), or else to an asymptotic notion of particle states within the interacting theory ([Bain, 2000]). More generally, [Ruetsche, 2011] gives an extended argument against pristine interpretation. While this difference in outlook is often implicit, D. Fraser (in response to [Wallace, 2001]) is transparently skeptical of even the cogency of an ontology with non-fundamental entities or which is not built on an exact similarity relation with the mathematical model at hand [Fraser, 2008, 857–8].
This difference in outlook clearly informs reactions to the particle problem. In addition to attitudes towards the ideal of pristine interpretation, there is a further extra-Haagian outlook operative under the surface of this literature. This outlook concerns the basic nature of the question at stake. For some, the question to be answered is circumscribed as a highly constrained puzzle within a formalization of our intuitions of particle-like things. Others, in contrast, aim to assess the viability of a particle interpretation understood much more loosely and flexibly from the outset. [Fraser, 2009a, 857] explains this difference (situating herself in the former camp) as coming down to solving a problem within a research program versus setting out on a different program altogether: those who, like Wallace and Bain (and in a different way [Halvorson and Clifton, 2002] and [Feintzeig et al., 2021]), aim at rehabilitating (our talk of) particles under a revised, emergent, or approximate notion of particle offer “a program, not a solution” to the original puzzle.
@@
For pristine interpreters, Haag’s theorem is a diagnostic no-go result concerning interpretation: QFT cannot be interpreted as a theory of particles. For instance [Halvorson and Clifton, 2002, 23] appeal to Haag’s theorem, among other no-go results, when they conclude that the theory of quantum fields “does not permit an ontology of localizable particles; and so, strictly speaking, our talk about localizable particles is a fiction”, useful though this fiction may be. As [Halvorson and Clifton, 2002, 20] briefly describe, Haag’s theorem precludes a particle ontology as follows. One first requires that there is a representation of the Weyl relations having a global occupation number operator. These are usually furnished with the interpretation of counting the number of particles in a state (i.e., it is the particle representation discussed in 2.1). [^22] See [Fraser, 2008, 845-846] for an exposition as to why these operators at apt for counting particles, rather than counting energy levels or other discrete quantum quantities. Since number operators only exist for representations unitarily equivalent to a free-field Fock representation (theorem 3.3 of [Chaiken, 1968]), and since Haag’s theorem implies that the latter cannot represent an interacting field, interacting field theory does not admit of a particle interpretation.
@@
[Fraser, 2008, 847-849] demonstrates the force of Haag’s theorem against particle interpretation in more detail. She describes three distinct attempts to generalize a quanta interpretation of free fields to interacting fields, arguing that each of these fail. The first of these attempts is foiled by Haag’s theorem; the others fail for different reasons. [^23] But they, like the approach in [Bain, 2013] to the particle problem, count the evasion of Haag’s theorem among their merits. The first attempt tries to rather straightforwardly extend the notion of particle available to free field theories to interacting ones, just as [Halvorson and Clifton, 2002] discuss. In the free case, we obtain a well defined number operator, whose eigenvectors may be interpreted as counting definite numbers of particles so long as we are in a Fock representation. But Haag’s theorem shows us that we cannot use a Fock representation of free fields to represent interacting ones. [^24] See also [Heathcote, 1989] pp. 91-97 for a discussion of the consequences of Haag’s theorem for Fock space representations and axiomatic approaches to modeling interactions, and see [Earman and Fraser, 2006] section 7 for their critical assessment thereof. And of course, when we are looking for a particle interpretation of QFT, we care significantly more about the interacting case than we do the free case.

### 06 Apr 2024 v3

@@ 4 The Framework: Assessment, Repair, and Renovation and Maintenance
Agreement (2) belies the different expectations one might have for the form of QFT. Beyond furnishing a method of generating predictions, one pre-Haagian appeal of the interaction picture was its intuitive representation of interaction dynamics. The asymptotic nature of calculations aside, interactions looked familiar: the full state space was time-evolved, from beginning to end, by a specified dynamical operator. As the agreement on (2) reflects, Haag’s theorem undermines the interaction picture’s ability to represent interactions in this way. And while renormalization delivers the agreed, reliable predictions of (1), it does so at the cost of further complicating attempts to represent interactions as originally hoped. But whether any of this is problematic, and if so to what degree, depends on one’s hopes and expectations. What should count as a fully satisfactory theory? as a fully satisfactory theory of relativistic QFT?
In thus organizing the various responses to Haag’s theorem, and in bringing the background commitments of the extra-Haagian outlooks into broad daylight, the framework itself does not adjudicate among the responses to Haag’s theorem. Rather, it clarifies the various lines of genuine debate, both at the level of disagreement of technical matters of fact, and at the level of deeper methodological commitments. As such, these outlooks are directly implicated in 5. Since there is no single “wner’ of QFT, let alone the edifice of science, the information that any inquirer of QFT will want—be they physicist, philosopher, or mathematician—will vary according to their own outlook. In our role as ‘general contractors’, we indulge these owners by providing sample advice that either arranges the responses according to their outlook and even provisionally adopts an outlook.
A final note before we proceed. While the responses to Haag’s theorem we review below proceed as if it presents a single problem demanding a singular response, this is not universal. In particular, [Ruetsche, 2011, Chs. 9–11] notes several potential problems that unitary inequivalence phenomena, Haag’s theorem among them, pose to a fundamental particle interpretation. Her response to these is pluralist insofar as she recognizes that “different interpretations will be indexed to—and adulterated by—different aspirations,” where these aspirations “adulterate because they arise along with particular applications of the theory” [Ruetsche, 2011, 246]. She argues that to demand a univocal interpretation, which would require a singular response to Haag’s theorem, would “stym[ie] attempts to harness the theory […] in fruitful projects of explanation and expansion” [Ruetsche, 2011, 260]. While this pluralist response is rich and worthy of discussion, we do not review it here because of its complexity and because it is not centered on Haag’s theorem per se.
@@
Abandon the IP; adopt Haag-Ruelle scattering
Interpret and assess formal variations of QFT
Theory Interpreters
In general no number operator for interacting QFT
Develop alternative ontology for QFT
Develop rigorous interacting QFT
@@
Renormalization defuses Haag’s theorem by breaking Poincaré invariance
Maintain current use of renormalization and regularization techniques
@@
Revise Wightman axioms to introduce new evolution parameter
@@
D. Fraser’s dissertation, Haag’s Theorem and the Interpretation of Quantum Field Theories with Interactions, set the stage for much of the philosophical literature on Haag’s theorem over the last sixteen years ([Fraser, 2006]). We will therefore give a more detailed assessment of this source than subsequent works. She writes, “Haag’s theorem frames answers to the question ‘What is QFT?’ ” (p. 171). There are several contenders for the prize of being QFT. Fraser divides what we have called canonical QFT into three subcategories: unrenormalized canonical QFT, renormalized canonical QFT, and canonical QFT with cutoffs. Contending against these for the prize of being QFT are several different formal variants of QFT: bottom up axiomatic work based on either the Wightman or Haag-Kastler axioms; constructive QFT; and algebraic QFT. Haag’s theorem shows us that unrenormalized canonical QFT is predicated on an inconsistent set of initial assumptions, and is therefore disqualified. Neither renormalization nor the introduction of cut-offs are satisfactory in the face of this inconsistency:
@@
[C]onsistency is also a relevant criterion because quantum field theory is, by definition, the theory that integrates quantum theory and the special theory of relativity. Consistency is relevant to QFT for theoretical reasons—not for practical reasons (e.g., the derivation of predictions). As a result, it is necessary to either formulate a consistent theory or else show that this criterion cannot be satisfied (i.e., that there is no consistent theory with both quantum and special relativistic principles). …The formal variant is the only variant that satisfies the criterion; its set of theoretical principles are both consistent and well motivated. [Fraser, 2009] (p. 563)
Thus we have the key dialectical currents of the Fraser-Wallace debate fueling this line of response to Haag’s theorem. Bedrock issues about what QFT even is, and what it takes to formulate it in a way that is apt for deep philosophical work on its conceptual-nomological machinery place Haag’s theorem in the spotlight: this no-go theorem regiments answers to the central question at hand, namely, What is QFT? In contrast, the typical theoretical particle physicist contents themselves with combining SR and QM to the best of their current ability, rather than demanding perfectly coherent and mathematically rigorous unification up front.
@@
Arguably, the best of their ability has improved over the decades such that, the average theoretical particle physicist could claim, we will inevitably find such a complete and honest marriage if one is forthcoming. In the meantime, we have successful, if incomplete, marriages of SR with QM with which we can keep experimental and theoretical particle physics going. While it is likewise a live option for D. Fraser (and presumably Earman) that an honest and complete marriage of QFT is not possible, this is a question of logic, not practical or empirical feasibility.141414 Of course, D. Fraser’s full corpus on the philosophy of QFT extends beyond the scope of issues in this present paper. In her dissertation she seemed optimistic that, with more time, the as-of-yet unfinished project of constructing a fully rigorous interacting QFT in four spacetime dimensions could be completed. Crucially, she allowed for the possibility that this will require revision to then-extant axiom systems ([Fraser, 2006] section 5). Subsequently, she put it to the proponents of renormalized and cutoff variants to wrestle with the question of whether such a completion is impossible: “Neither the infinitely renormalized nor cutoff variant furnishes an argument that a consistent formulation of QFT is impossible; such an argument would require making the case that the axiomatic program cannot be completed.” [Fraser, 2009] (p. 563). See [Koberinski and Fraser, 2023] for her most recent views on the epistemological status of effective field theory approaches in particle physics.
@@
We note briefly here that the physicists’ attitude of settling for good-enough-to-be-getting-on-with variants of QFT clearly extends far beyond Haag’s theorem. As just one other example, theoretical particle physicists are not bothered when they learn that their QFT path integrals cannot actually be integrals, given that there is no measure on the infinite-dimensional space of paths. Sorting out where path integrals properly live in the mathematical universe is a mathematician’s job—not a physicist’s. Meanwhile, that these path integrals are currently mathematically homeless in no way diminishes their calculational and conceptual fruitfulness for theoretical physics. Haag’s theorem is one result among many indicating that the formal and canonical variants of QFT are not communicating well with each other: either, as Earman and D. Fraser see it, the foundations of QFT are in excellent shape, and they are waiting for more mathematical success in erecting interacting models of realistic matter in a realistic number of spacetime dimensions; or, as Wallace sees it, QFT needs no foundational work since it does not really even need (mathematically) rigorous foundations. We, the authors, think that this communication breakdown between formal and informal variants of QFT is problematic. We will return to these concerns in section 5.
@@
In addition, Earman and D. Fraser recognize that an explanation is needed for the success of the interaction picture and perturbation theory in the face of Haag’s theorem. They “suspect that the full explanation will have a number of different pieces” and offer their analysis of Haag-Ruelle scattering as one such piece (p. 322). The Duncan and Miller response discussed below (section 4.3) should be thought of as adding one additional such piece, not refuting Earman and D. Fraser.
@@
As a consequence of this logical diagnosis and logical remedy, Earman and D. Fraser advocate for future philosophical work and resources to be deployed in philosophical projects about formal variants of QFT and their attendant non-interaction picture approaches to interactions. There are at least two such broad philosophical projects that benefit from the ways in which practicing mathematicians have “digested the lesson of Haag’s theorem” [Earman and Fraser, 2006] (p. 334). First there is the project of interpreting the mathematical structures used in algebraic QFT and in constructive QFT. Second, and relatedly, there is the subtle matter of assessing the implications of theorems of axiomatic approaches to QFT for the philosophy of those areas of fundamental physics that make use of QFT.
@@ 4.2 Theory Interpreters—The Particle Problem
There is a segment of the philosophical literature addressing Haag’s theorem that is primarily motivated by the question, as put by Ruetsche, “Is particle physics particle physics?” [Ruetsche, 2011, 190]. The issue is whether or not the quantum field theories employed in particle physics admit of particle interpretations. The lines of argument for and against particle interpretations are numerous, and we direct interested readers to [Ruetsche, 2011] ch. 9, and [Fraser, 2021] as entry points to these debates. Here we focus on the hurdle that Haag’s theorem poses for particle interpretations.
[Fraser, 2008] advances the argument from Haag’s theorem to the elimination of particles from our ontology. ([Halvorson and Clifton, 2002, 20] briefly sketch the initial segment of this argument.) The heart of the argument runs, roughly, as follows. At a minimum, a particle interpretation requires that the theory admits of a formulation in which a number operator exists—you cannot very well say that the theory says it is possible that matter is fundamentally particles if the theory also says that it is not possible (in theoretical principle) for those particles to be counted. The Fock space representation of a free field yields a well-defined number operator as well as the correct relativistic energies suitable for a particle interpretation. Might we then, by means of the IP, use the Fock space of a free field for representing and counting particle states in an interacting system as well? Haag’s theorem blocks this path, since it entails that representations of the ETCCRs for free and interacting systems are unitarily inequivalent. For this and other reasons (see Assessment below), Fraser, Halvorson, and Clifton conclude that a particle interpretation for QFT is a no-go, and so there are no particles. Others ([Wallace, 2011] and [Bain, 2011]) disagree and seek to rehabilitate an emergent particle interpretation despite Haag’s theorem.
@@
This sub-literature is unified by a decidedly interpretive task: the authors aim (in this context) to answer questions about what ontological commitments are supported by QFT. Thus, this literature is not ultimately motivated by considerations of theory consistency, application, development and innovation, or conceptual-nomological machinery (“how does this theory work?”) that guide the work of groups surveyed in other subsections. Although such considerations sometime arise in this literature (as, indeed, matters of theory interpretation sometimes arises as a secondary consideration in other segments of the Haag’s theorem literature), they do so as steps within larger arguments ultimately aimed at interpretation. (And of course, many authors take up distinctive projects across this interpretation/not-interpretation divide.)
There is a further extra-Haagian outlook operative under the surface of this literature. This outlook concerns the basic nature of the question at stake. For some, the question to be answered is circumscribed as a highly-constrained puzzle: given QFT in a mathematically rigorous form, is it possible to construct a number operator for interacting QFT? Others, in contrast, aim to assess the viability of a particle interpretation understood much more loosely and flexibly from the outset. [Fraser, 2009, 857] explains this difference (situating herself in the former camp) as coming down to solving a problem within a research program versus setting out on a different program altogether: those who, like Wallace and Bain (and in a different way [Feintzeig et al., 2021]), aim at rehabilitating (our talk of) particles under a revised, emergent, or approximate notion of particle offer “a program, not a solution” to the original puzzle. Ruetsche agrees with Fraser’s analysis of the way in which Haag’s theorem precludes a particle interpretation of interacting QFT writ large. However, Ruetsche argues for a multifaceted, less “pristine” approach to theory interpretation. For Ruetsche, the argument from Haag’s against particles applies in certain context, but not in others, such that “[s]ometimes particle physics is, adulteratedly, particle physics, and that’s a good thing” [Ruetsche, 2011, 260].
@@
Thus, there are extra-Haagian disagreements over the method of theory interpretation. There is also an extra-Haagian disagreement over the standards of theory formulation required before the interpretive task can begin. For [Halvorson and Clifton, 2002] and [Fraser, 2008], the theory to be interpreted is the formal, mathematically well-posed version of the scientific achievement under study ([Fraser, 2009], [Fraser, 2011]). No-go theorems proved within these formal theoretical structures, then, powerfully and persuasively constrain interpretive options  [^18] See [Freeborn et al., 2023] for further discussion on alternative stances toward no-go theorems in general, and Haag’s theorem in particular. [Fraser, 2008] is explicit on this point. In describing the role of Haag’s theorem for blocking one attempt at developing a particle interpretation of interacting QFT, she writes:
@@
Moreover, according to Wightman, it is a “necessary condition” that any physically sensible interacting field theory use a representation of the ETCCRs that is unitarily inequivalent to a Fock representation.Thus, Haag’s theorem poses a problem for a particle interpretation provided that one is committed to mathematical rigor in theory formulation and interpretation.  [^19] See [Hans Halvorson, 2007] p. 731-732 for a concise statement of this attitude towards interpretation and mathematical precision in theory formulation.
In contrast, [Bain, 2000] and [Wallace, 2001] take a more flexible and opportunistic approach to the question of particle interpretations for QFT. To them, Wightman’s “necessary condition” looks more like an interesting suggestion or a useful starting point. From such a starting point, they turn their attention toward opportunities within the messy details of how interacting QFT gets applied to develop a serviceable notion of particles without meeting the stringent requirements of finding a representation unitarily equivalent to a Fock space representation. The usual response along these lines appeals to a scale-variant or emergent particle interpretation ([Wallace, 2011, 124], [Wallace, 2001])202020[Wallace, 2001] aims to give a concrete treatment of the emergence of particle structure from quantum fields. He aims to explain how the particle concept comes to be useful for interpreting a theory that seems prima facie to be about very un-particle-like entities known as fields. This is not a solution to the highly-constrained puzzle (constructing a number operator for interacting QFT) that [Fraser, 2008] and [Halvorson and Clifton, 2002] aim to address., or else to an asymptotic notion of particle states within the interacting theory ([Bain, 2000]). While this difference in outlook is often implicit, D. Fraser (in response to [Wallace, 2001]) is transparently skeptical of even the cogency of an ontology with non-fundamental entities or which is not built on an exact similarity relation with the mathematical model at hand [Fraser, 2008, 857–8].
@@
For [Fraser, 2008] and [Halvorson and Clifton, 2002], Haag’s theorem is a diagnostic no-go result concerning interpretation: QFT cannot be interpreted as a theory of particles. For instance [Halvorson and Clifton, 2002, 23] appeal to Haag’s theorem, among other no-go results, when they conclude that the theory of quantum fields “does not permit an ontology of localizable particles; and so, strictly speaking, our talk about localizable particles is a fiction”, useful though this fiction may be. As [Halvorson and Clifton, 2002, 20] briefly describe, Haag’s theorem precludes a particle ontology as follows. One first requires that there is a representation of the Weyl relations having a global occupation number operator. These are usually furnished with the interpretation of counting the number of particles in a state (i.e., it is the particle representation discussed in 2.1). [^21] See [Fraser, 2008, 845-846] for an exposition as to why these operators at apt for counting particles, rather than counting energy levels or other discrete quantum quantities. Since number operators only exist for representations unitarily equivalent to a free-field Fock representation (theorem 3.3 of [Chaiken, 1968]), and since Haag’s theorem implies that the latter cannot represent an interacting field, interacting field theory does not admit of a particle interpretation.
@@
[Fraser, 2008, 847-849] demonstrates the force of Haag’s theorem against particle interpretation in considerably more detail. She describes three distinct attempts to generalize a quanta interpretation of free fields to interacting fields, arguing that each of these fail. The first of these attempts is foiled by Haag’s theorem; the others fail for different reasons. [^22] But they, like the approach in [Bain, 2013] to the particle problem, count the evasion of Haag’s theorem among their merits. The first attempt tries to rather straightforwardly extend the notion of particle available to free field theories to interacting ones, just as [Halvorson and Clifton, 2002] discuss. In the free case, we obtain a well defined number operator, whose eigenvectors may be interpreted as counting definite numbers of particles so long as we are in a Fock representation. But Haag’s theorem shows us that we cannot use a Fock representation of free fields to represent interacting ones. [^23] See also [Heathcote, 1989] pp. 91-97 for a discussion of the consequences of Haag’s theorem for Fock space representations and axiomatic approaches to modeling interactions, and see [Earman and Fraser, 2006] section 7 for their critical assessment thereof. And of course, when we are looking for a particle interpretation of QFT, we care significantly more about the interacting case than we do the free case. Given this argument [Halvorson and Clifton, 2002] and [Fraser, 2008] conclude that the fundamental ontology of the actual world does not include any particles. In contrast, [Bain, 2013], [Wallace, 2001], and [Feintzeig et al., 2021] aim (in their own distinctive ways) to recover a non-fundamental, emergent notion of particles.